Find the quadratic function y=ax2+bx+c whose graph passes through the given points.
​(3,−11​), ​(−2​,−16​), ​(1,−1​)

Answer by stanbon(75887)

Find the quadratic function y=ax2+bx+c whose graph passes through the given points.
​(3,−11​), ​(−2​,−16​), ​(1,−1​)
-----
(3,-11) gives:: 9a + 3b + c = -11
(-2,-16) gives: 4a - 2b + c = -16
(1,-1) gives:: a + b + c = -1
--------
Use any method you know to solve the system for a,b,c.
-----------------
a = -2 ; b = 3; c = -2
-----
Equation:
y = -x^2 + 3x -2
--------------
Cheers,
Stan H.
------------